Calculate for each graph the elasticity of demand. Then, evaluate if unit elastic, inelastic or elastic, why?

A) B) C) 17 09 34 13 1a 130 38 137   
Answer #1

Answer to graph A : Elasticity of demand = - 0.46 or 0.46 ( absolute value) and it is inelastic

Explanation:

Ed = \Delta Q /\Delta P * [(P1 + P2) / (Q1 +Q2)]

= {( 130 -112 ) / ( 21 - 29 ) } * { ( 29 +21 ) / ( 112 +130)

= (18 / - 8 ) * ( 50 / 242 )

= - 2.25 * 0.2066

= - 0.46

Answer to graph B : Elasticity of demand = - 4.96 or 4.96 ( absolute value) and it is elastic

Explanation:

Ed = \Delta Q /\Delta P * [(P1 + P2) / (Q1 +Q2)]

= {( 38 -187 ) / ( 17 - 13 ) } * { ( 13 +17 ) / ( 187 + 38 )

= ( - 149 / 4 ) * ( 30 / 225 )

= - 37.25 * 0.1333

= - 4.96

Answer to graph C : Elasticity of demand = - 0.0894 or 0.0894 ( absolute value) and it is inelastic

Explanation:

Ed = \Delta Q /\Delta P * [(P1 + P2) / (Q1 +Q2)]

= {( 65 - 69 ) / ( 34 - 17 ) } * { ( 17 +34 ) / ( 69 + 65 )

= ( - 4 / 17 ) * ( 51 / 134 )

= - 0.2352 * 0.3805

= - 0.0894

If Ed is greater than 1 , then the demand is elastic.

If Ed is less than 1 , then the demand is inelastic.